OpenStudy (anonymous):

\[d/dx \sum_{o}^{\inf} [(x^(2n)))/n!]\] I will award a medal if you help me get the right answer!!

OpenStudy (anonymous):

\[d/dx \sum_{n=o}^{\inf} [(x^(2n)))/n!]\]

OpenStudy (accessdenied):

Derivative of a sum is just the derivative of the individual terms, right? So you can interchange the sum and derivative because we are essentially taking: \( \dfrac{d}{dx} \left( f_1 (x) + f_2 (x) + f_3 (x) + \cdots \right) = \dfrac{d}{dx} f_1 (x) + \dfrac{d}{dx} f_2 (x) + \dfrac{d}{dx} f_3(x) + \cdots \)

OpenStudy (anonymous):

okay so that leaves us with what?

OpenStudy (accessdenied):

\( \displaystyle \dfrac{d}{dx} \sum_{n=0}^{\infty} x^{2n} / n! = \sum_{n=0}^{\infty} \dfrac{d}{dx} \left( x^{2n} / n! \right) \)

OpenStudy (anonymous):

okay

OpenStudy (accessdenied):

Then take the derivative, using the power rule. d/dx(x^a) = a x^(a-1). d/dx (x^(2n)) = ?

OpenStudy (anonymous):

2nx^(2n-1)

OpenStudy (accessdenied):

Yep, looks good. So we would now have: \( \displaystyle = \sum_{n=0}^{\infty} 2n x^{2n - 1} / n! \) There may be some minor simplifications (n=0 gives us 0, so it can be lost. n and n! = n(n-1)! have common factors to cancel), but that may be done as you feel necessary here.

OpenStudy (anonymous):

okay so how do we find what it equals?

OpenStudy (accessdenied):

Like, we want to find the actual function with that series expansion?

OpenStudy (anonymous):

the answer options are as follows a) -sin(x^2) b) -2xsin(x^2) c) cos(x^2) d) \[e^x^2 \]

OpenStudy (anonymous):

or e) 2xe^(x^2)

OpenStudy (anonymous):

d) was supposed to say e^(x^2)

OpenStudy (anonymous):

any idea?

OpenStudy (accessdenied):

Hm.. if you know the expansion for e^x: \( \displaystyle e^{x} = \sum_{n=0}^{\infty} \dfrac{x^n}{n!} \) which when we have \(x \to x^2\), this becomes what we started with: \( \displaystyle e^{x^2} = \sum_{n=0}^{\infty} \dfrac{x^{2n}}{n!} \) When we take this function's derivative (which is equivalent to above, this is d/dx(x^2) e^(x^2) = 2x e^(x^2). Otherwise we can put the derivative expression in the same terms as the e^(x^2) expansion by making those simplifications I mentioned above: \( \displaystyle \sum_{n=0}^{\infty} 2n \dfrac{x^{2n-1}}{n!} = \sum_{n=1}^{\infty} 2 \dfrac{x^{2(n-1)+1}}{(n-1)!} = 2x \color{blue}{\sum_{n=0}^{\infty} \dfrac{x^{2n}}{n!}} \)

OpenStudy (anonymous):

okay, so do we get d as the answer then?

OpenStudy (accessdenied):

I condensed a lot of steps in that last equation, but if anything is unclear I can elaborate that.

OpenStudy (anonymous):

No, i understand the steps you took!

OpenStudy (accessdenied):

(d) is just e^(x^2), but that was just the original thing we started with. What we took was the derivative of the e^(x^2) expansion, so it is the equivalent of taking the derivative of e^(x^2).

OpenStudy (anonymous):

so would it be e then?

OpenStudy (accessdenied):

d/dx ( e^(x^2) ) is just a chain rule application, we take the derivative of the x^2 times the original function (because exponential). 2x * e^(x^2), would be (e) yes.

OpenStudy (anonymous):

Okay, thanks so much for your help!

OpenStudy (accessdenied):

Yup, glad to help! :)